OpenStudy (anonymous):

Plz explain 23-4x=27 I know the answer is -4 just learning this , this week. Thanks

OpenStudy (the_fizicx99):

23 - 4x = 27 You want to isolate x, and what that basically means you want to have it alone on one side, Subtract 23 from both sides, and divide both sides by -4

OpenStudy (anonymous):

Can you only ask one math question At a time?

OpenStudy (the_fizicx99):

Sort off, it's one per thread, a question or doubt about this problem is fine

OpenStudy (anonymous):

Ok so 27-23 and 23. -23. ? 27-23= ~4

OpenStudy (the_fizicx99):

27 - 23 = 4 So now you have -4x = 4 Divide both sides by -4, you get?

OpenStudy (anonymous):

23 - 4x = 27 23 - 27 - 4x = 27 - 27 -4 - 4x = 0 -4 + 4 -4x = 0 + 4 -4x = 4 -4x / -4 = 4 / -4 Therefore, x = -1

OpenStudy (anonymous):

I meant i know the answer os -1, but its still confusing, i thought i subtracció 23 from 23 and 27 first

OpenStudy (the_fizicx99):

You want it to be explained in Spanish, maybe you understand it better?

OpenStudy (anonymous):

No im not spanish.

OpenStudy (the_fizicx99):

ah k, I don't know what "subtracció " means...

OpenStudy (anonymous):

I think Tralaman mistyped somehow from subtract to subtraccio

OpenStudy (anonymous):

Sorry , i meant subtracted

OpenStudy (the_fizicx99):

no you subtract 23 from both sides, 23-23 is 0, 27-23 is 4 now your last step is to divide by * to isolate x

OpenStudy (anonymous):

Thanks for explaining step one . The answer is 4. Then divide 4 into 23 and 27?

OpenStudy (anonymous):

4x/4 and 4/4

OpenStudy (the_fizicx99):

No no,